OpenStudy (anonymous):

Choose the expression that completes the equation. 33 ÷ 3 + 10 = 2(10 + 1) 10(1 + 1) 3(6 + 1) 5(3 + 5)

OpenStudy (anonymous):

What is 33/3 +10?

OpenStudy (anonymous):

3/3=11+10= 21 now, find which of those choices also equals 21 (:

OpenStudy (anonymous):

33 divided by 3 plus 10

OpenStudy (anonymous):

3(6 + 1) which equals 21. Do you still need help? xD

OpenStudy (anonymous):

2(10+1)!

OpenStudy (anonymous):

@Yvonne_979 that equals 22, not 21

OpenStudy (anonymous):

my bad ! lol.